# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

### v2

@@ 8.2.2. FedFDP
Heterogeneity. Table 5 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $Dir(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
@@
Figure 3. Impact of different $\lambda$ in FedFDP algorithm
[…]
Figure 4. Impact of different $C$ in FedFDP algorithm
[…]
Impact of Different $\lambda$. We investigated the correlation between $\lambda$ and fairness $\Psi$ in the FedFDP algorithm. As depicted in Figure 3, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
Impact of Different $C$. In Section 5.1, we emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
@@
9. Conclusion
@@ Appendix A Appendix
A.1. Proof of Theorem 1
@@
(28) $\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
@@
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich, 2018). Therefore,
@@
(29) $\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
A.1.1. Key Lemma
Lemma 0.
@@ A.1.2. Proof of Lemma 1
(30) $\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$
@@
A.2. Proof of Theorem 1
$proof.$ We will prove Theorem 1 in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 0.
(RDP privacy budget of SGM(Mironov et al., 2019)). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
(31) $R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$
@@ A.2. Proof of Theorem 1
(32) $\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$
@@
Finally, the existing work (Mironov et al., 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
(33) $A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$
Definition 0.
(Composition of RDP(Mironov, 2017)). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$

### v3

@@ 8.1 Results
Heterogeneity. Table 2 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
@@
Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
Figure 4: Impact of different $C$ in FedFDP algorithm.
Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
Impact of Different $\lambda$. We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm. As depicted in Figure 3, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
Impact of Different $C$. We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
@@
Impact of Different $\sigma$. We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $[1,1.5,2,2.5,3]$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T=[6,115,268,463,708]$ rounds of communication. As shown in Figure 5, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence. When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value. As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
9 Conclusion
@@ Appendix A Appendix
$F(\mathbf{w}$) Global loss function
$F_{i}(\mathbf{w}$) Local loss function for client $i$
$T$ Communication rounds
$t$ Communication round index
$D_{i}$ Dataset of client $i$
$D$ The union of $D_{i}$
$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
$\eta$ Gradient descent step size
$\lambda$ Fairness parameter
$N$ Total number of clients
$q$ batch sample ratio
$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
$\sigma$ Noise multiplier for gradient
$\sigma_{l}$ Noise multiplier for loss
$C$ Original clipping bound for gradient
$C_{l}$ Original Clipping bound for loss
|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
Table 4: Summary of main notations
A.1 Extra Experiment
We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running Windows 10.
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
Table 5: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
Table 6: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Default Settings
We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al. 1998), FashionMNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR10 (Krizhevsky 2009). And we comparised our algorithms to six baselines:
  1. 1.
    FedAvg(McMahan et al. 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
  2. 2.
    SCAFFLOD(Karimireddy et al. 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
  3. 3.
    FedProx(Li et al. 2020a): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  4. 4.
    FedDyn(Acar et al. 2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  5. 5.
    ALI-DPFL(Ling et al. 2024): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
  6. 6.
    q-FFL(Li et al. 2020b): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
For all datasets, we adopted a widely used heterogeneous settings (Lin et al. 2020; Li, He, and Song 2021) to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ (Lin et al. 2020; Wang et al. 2020). We use a 4-layer CNN architecture (McMahan et al. 2017) which consists of two convolutional layers and two fully connected layers as model architecture.
Extended description of FedFair
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
14
15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
16       for batch $b\in\mathcal{B}_{i}$ do
17             $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
18             $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
20
22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
23
Algorithm 2 FedFair
  • Server (lines 2-2): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 2-2): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 2-2. For line 2, we execute gradient descent by using $\eta_{i}$.
Result of FedFair
Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich 2003). Performance and Fairness. Table 5 presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
Heterogeneity. Table 6 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
A.2 Proof of Theorem 1
@@
$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (17)
@@
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich 2018). Therefore,
@@
$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (18)
Key Lemma
Lemma 1
@@
$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (19)
@@
A.3 Proof of Theorem 2
When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$. Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (14) to obtain a function in terms of $\lambda$,
$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (20)
$\mathcal{F}^{\prime}(\lambda)=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$ (21)
Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
$\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$ (22)
$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
$=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$ (23)
The derivative of $\mathcal{G}(\lambda)$ is given by
$\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$ (24)
It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
$\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$ (25)
@@ A.3 Proof of Theorem 2
  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
[…]
A.4 Proof of Theorem 3
As Algorithm 1 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 1-1 and lines 1-1 of the Algorithm 1, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
Privacy Loss of $\mathbf{w}_{t+1}^{i}$.
Theorem 4
$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (26)
where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 3
(RDP privacy budget of SGM(Mironov, Talwar, and Zhang 2019)). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (27)
where
$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (28)
@@ A.4 Proof of Theorem 3
Finally, the existing work (Mironov, Talwar, and Zhang 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (29)
Definition 4
(Composition of RDP(Mironov 2017)). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
@@
Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
Theorem 5
$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (30)
The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
Proof Loss of FedFDP.
Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 2 to convert it to $(\epsilon,\delta)$-DP. Therefore, the Theorem 3 is hold.
Lemma 2
(Conversion from RDP to DP (Balle et al. 2020)). if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.